Write the equation of the given parabola in general form:
-14x + 2y^2 -8y =20

Answer by stanbon(75887)

Write the equation of the given parabola in general form:
-14x + 2y^2 -8y =20
---------------
Rearrange:
14x+20 = 2y^2-8y
14x+20+8 = 2(y^2-4y+4)
14x + 28 = 2(y-2)^2
---
14x = 2(y-2)^2-28
---
x = (1/7)(y-2)^2-2
---
OR
7x+10 = y^2-4y
x = (1/7)y^2-(4/7)y-(10/7)
==============================
Cheers,
Stan H.